For this question, we consider a two-allele, one locus trait. Here, assume the trait
that this locus affects is coat thickness.
In a juvenile population of bison, they commonly experience a harsh winter. The AA
individuals lose about 50% of their number, the Aa individuals lose about 30%, and
the aa individuals lose about 90% of their number. Before the winter, there are
1000 AA individual, 300 Aa individuals, and 700 aa individuals (generation 1).
a) What are the relative fitnesses of each genotype i.e. (WAA, Waa)?


b) What are the expected allele frequencies in the next generation (assume
N=1000 for generation 2)?


c) Draw an approximate graph illustrating the change in frequency of the A and
a alleles over time in a different population where the starting frequency of
the a allele =0.2. Explain your answer i.e. why are your curves shaped the
way they are?


d) Do you expect the genetic variability in this population to be maintained or
lost over time? Why?
Answer #1

(a). Relative fitness of a genotype is given by the survival and/or reproductive rate of a genotype (or phenotype) relative to the maximum survival and/or reproductive rate of other genotypes in the population.

The survival percentage of AA is 50%, Aa is 70% and of aa is 10%

WAA- 50/70 =0.714

WAa= 70/70 = 1

Waa= 10/70= 0.142

(b) Relative frequency of allele is given by hardy weinberg principle-Date am 2 Arnd -According auelie pen vemeie on gemelatlom to teu mert No , a = 140 pepelakon 出 0,775 IS6o oultelu ADate is60 auela ト. 2o.77